# Exotic synchronization in continuous time crystals outside the symmetric subspace

### v1

@@ Abstract
Here, $\mathcal{L}_{\Omega_{\alpha},\kappa_{\alpha}}(\rho)$ describes the dynamics of uncoupled CTCs, and $\Gamma_{\alpha,\beta}$ is the coupling strength between two CTCs $\alpha$ and $\beta$. Such an interaction resulted in the observation of the seeding effect and the synchronization of detuned CTCs considering only the symmetric subspace Hajdušek et al. (2022). We now also take into account asymmetric subspaces and assume that all the CTCs are identical, $\mathcal{L}_{\Omega_{\alpha},\kappa_{\alpha}}=\mathcal{L}_{\Omega,\kappa},~{}\forall\alpha$. Even if they are identical, different choices of initial states can lead to surprising disparate effects in the dynamics. We investigate the mean-field dynamics of $n=20$ coupled CTCs in Fig. 3, where all CTCs are initialized in asymmetric subspaces belonging to different values of $m$. The corresponding frequency and the dynamics of $m_{z}^{\alpha}$ of uncoupled CTCs depending on the value $m$ are shown in Fig. 3(a) and Fig. 3(d), respectively. Depending on the frequency distribution given by Fig. 3(a), the system splits into two ensembles ${\varepsilon_{1}}$ and ${\varepsilon_{2}}$ which are indicated by black and red colors, respectively. First, we consider the case where CTCs within the same ensemble are coupled with a finite interaction strength but there is no coupling between CTCs of different ensembles such that $\Gamma_{\alpha,\beta}=\Gamma\delta_{\varepsilon_{i}\varepsilon_{j}}~{}\forall~{}(\alpha\in\varepsilon_{i},\beta\in\varepsilon_{j})$. This results in the synchronization of CTCs within the same ensemble as shown by Fig. 3(b,e). Surprisingly, considering all-to-all coupled CTCs ($\Gamma_{\alpha,\beta}/\kappa=\Gamma/\kappa=0.35~{}\forall~{}(\alpha,\beta)$) leads to the synchronization of the CTCs only in one ensemble (here $\varepsilon_{1}$), while the CTCs in the other ensemble (here $\varepsilon_{2}$) exhibit unsynchronized dynamics as shown in Fig. 3(c,f). Such surprising patterns of partial synchronization in systems of coupled identical non-linear oscillators were first observed in Kuramoto and Battogtokh (2002) and were called chimera states Abrams and Strogatz (2004); Zakharova (2020). First strides towards chimera states in quantum systems involved exploring one-dimensional arrays of quantum oscillators Bastidas et al. (2015) and periodically driven networks of spins Sakurai et al. (2021). Our work demonstrates a chimera state in a system of coupled identical CTCs. Here, the chimera state results from including the asymmetric subspaces into the dynamics of identical CTCs, allowing us to choose an initial state with different frequencies in different subspaces Ujjwal et al. (2017). Increasing the coupling strength results in the synchronization of all CTCs (see Appendix B). Another example of an exotic synchronization regime that can be realized in coupled identical CTCs is cluster synchronization where the system separates into two differently synchronized clusters. Note that such complex dynamics can be obtained for arbitrary initial conditions (see Appendix C).
Conclusions.—
@@
In this work, we focused on the fundamental question of whether continuous time crystals in spin models only exist if initialized in the symmetric subspace. Our first main result is that the inclusion of asymmetric subspaces leads to time-crystal behavior even if the ratio of drive to dissipation strength $\Omega/\kappa$ is less than one. A second result is that taking into account asymmetric subspaces leads to multistability: depending on the parameter values, the system can exhibit a time crystal or melted phase for different initial states. For an ensemble of coupled CTCs, this leads to our third result, the existence of chimera states and cluster synchronization, which is not possible for CTCs realized in the symmetric subspace. Our results provide new insights into the dynamics of time-crystal phases and point out strategies in the design of quantum networks that may lead to the observation of exotic synchronization phenomena. An interesting future direction is to study the effect of weak permutational symmetry-breaking interactions, which allow the system to move between various spin sectors. Such interactions are inevitable in various experimental settings and will offer further valuable insights into the fundamental characteristics of time crystals.
@@ Acknowledgements.
P.S. and S.V. thank R. Fazio for discussions. Numerical calculations were performed using QuTiP Johansson et al. (2012, 2013) and PIQS Shammah et al. (2018). M.H. was supported by [MEXT Quantum Leap Flagship Program] Grants No. JPMXS0118067285 and No. JPMXS0120319794. P.S. and C.B. acknowledge financial support from the Swiss National Science Foundation individual grant (grant no. 200020 200481). S.V. acknowledges support from DST-QUEST grant number DST/ICPS/QuST/Theme-4/2019. While preparing the manuscript, we became aware of related work Iemini et al. (2023).
@@
Appendix A Appendix A: Mean-field equations for coupled continuous time crystals
@@
Here we discuss the mean-field dynamics of the coherently coupled CTCs described by Eq. (3) of the main text. These equations are formulated in terms of the rescaled spin operators $m_{\alpha}=\langle\hat{S}_{\alpha}\rangle/S$ and are defined as follows,
$\frac{d\langle\hat{m}^{x}_{\alpha}\rangle}{dt}$ $=\kappa\langle\hat{m}^{x}_{\alpha}\rangle\langle\hat{m}^{z}_{\alpha}\rangle+\frac{\Gamma}{n}\langle\hat{m}^{z}_{\alpha}\rangle\sum_{\beta\neq\alpha}\langle\hat{m}^{y}_{\beta}\rangle,$
$\frac{d\langle\hat{m}^{y}_{\alpha}\rangle}{dt}$ $=-\Omega_{\alpha}\langle\hat{m}^{z}_{\alpha}\rangle+\kappa\langle\hat{m}^{y}_{\alpha}\rangle\langle\hat{m}^{z}_{\alpha}\rangle-\frac{\Gamma}{n}\langle\hat{m}^{z}_{\alpha}\rangle\sum_{\beta\neq\alpha}\langle\hat{m}^{x}_{\beta}\rangle,$ (4)
@@
$\frac{d\langle\hat{m}^{z}_{\alpha}\rangle}{dt}$ $=\Omega_{\alpha}\langle\hat{m}^{y}_{\alpha}\rangle-\kappa\left[\langle\hat{m}^{x}_{\alpha}\rangle^{2}+\langle\hat{m}^{y}_{\alpha}\rangle^{2}\right]+\frac{\Gamma}{n}\left[\langle\hat{m}^{y}_{\alpha}\rangle\sum_{\beta\neq\alpha}\langle\hat{m}^{x}_{\beta}\rangle-\langle\hat{m}^{x}_{\alpha}\rangle\sum_{\beta\neq\alpha}\langle\hat{m}^{y}_{\beta}\rangle\right].$
@@
Appendix B Appendix B: Chimera state and synchronization via seeding
@@
The initial state leading to the chimera state shown in Fig. 3 of the main text is chosen such that the absolute value of corresponding eigenvalue $\{|\lambda_{1}^{\alpha}|;\alpha=1\ldots 20\}$ for the fixed point $\mathcal{M}_{1}$ has an equal distribution from $0.2$ to $0.25$ for CTCs in ensemble $\varepsilon_{1}$ (represented by black color) and from $0.75$ to $0.85$ for CTCs in ensemble $\varepsilon_{2}$ (represented by red color). For the given initial state, we investigate the effect of increasing the coupling strength between the all-to-all coupled CTCs on the synchronization properties. For coupling strength $\Gamma/\kappa=0.5$, all the CTCs in ensemble $\varepsilon_{1}$ are synchronized while CTCs in ensemble $\varepsilon_{2}$ only show partial synchronization, as shown in Figs. 4(a,d). Increasing $\Gamma$ we see more complex dynamics (not shown here), but the system continues to exhibit a chimera state. At $\Gamma/\kappa\approx 0.605$, a transition occurs: the strength of the coupling leads to the melting of the CTCs of ensemble $\varepsilon_{1}$ as depicted in Figs. 4(b,e). Further increasing the coupling strength leads to the synchronization of the CTCs in ensemble $\varepsilon_{2}$ which seeds oscillations to ensemble $\varepsilon_{1}$. As a result, at coupling strength $\Gamma/\kappa=1.2$, all the CTCs oscillate with a common frequency as shown in Figs. 4(c,f). Such a synchronization via seeding is consistent with what was observed in Hajdušek et al. (2022).
@@
Figure 4: Normalized Fourier transform $\mathcal{F}[m^{z}_{\alpha}]$(c-d) and time evolution of $m^{z}_{\alpha}$ (d-f) for coupling strength $\Gamma/\kappa=0.5$(a,d), $\Gamma/\kappa=0.605$(b,e) and $\Gamma/\kappa=1.2$(c,f).
Appendix C Appendix C: Cluster synchronization

### v2

@@ Abstract
Here, $\mathcal{L}_{\Omega_{\alpha},\kappa_{\alpha}}(\rho)$ describes the dynamics of uncoupled CTCs, and $\Gamma_{\alpha,\beta}$ is the coupling strength between two CTCs $\alpha$ and $\beta$. Such an interaction resulted in the observation of the seeding effect and the synchronization of detuned CTCs considering only the symmetric subspace Hajdušek et al. (2022). We now also take into account asymmetric subspaces and assume that all the CTCs are identical, $\mathcal{L}_{\Omega_{\alpha},\kappa_{\alpha}}=\mathcal{L}_{\Omega,\kappa},~{}\forall\alpha$. Even if they are identical, different choices of initial states can lead to surprising disparate effects in the dynamics. We investigate the mean-field dynamics of $n=20$ coupled CTCs in Fig. 3, where all CTCs are initialized in asymmetric subspaces belonging to different values of $m$. The corresponding frequency and the dynamics of $m_{z}^{\alpha}$ of uncoupled CTCs depending on the value $m$ are shown in Fig. 3(a) and Fig. 3(d), respectively. For simplicity, we consider two subgroups based on the separation between the frequency distribution of uncoupled CTCs, though a general distribution could result in more than two subgroups. In Fig. 3(a), the system is split into two ensembles, ${\varepsilon_{1}}$ (black) and ${\varepsilon_{2}}$ (red), with uniform frequency distributions: ${\varepsilon_{1}}$ ranges from 0.2 to 0.25, and ${\varepsilon_{2}}$ from 0.75 to 0.85.
We use Pearson’s correlation coefficient as an indicator of synchronization. For any two coupled CTCs $\alpha$ and $\beta$, it can be defined as follows
$\mathcal{C}_{\alpha\beta}=\frac{\langle{m^{z}_{\alpha}m^{z}_{\beta}}\rangle-\langle{m^{z}_{\alpha}}\rangle\langle{m^{z}_{\beta}}\rangle}{\sqrt{(\langle{m^{z}_{\alpha})^{2}}\rangle-\langle{m^{z}_{\alpha}}\rangle^{2})(\langle{(m^{z}_{\beta})^{2}}\rangle-\langle{m^{z}_{\beta}}\rangle^{2})}},$ (4)
[…]
First, we consider the case where CTCs within the same ensemble are coupled with a finite interaction strength but there is no coupling between CTCs of different ensembles such that $\Gamma_{\alpha,\beta}=\Gamma\delta_{\varepsilon_{i}\varepsilon_{j}}~{}\forall~{}(\alpha\in\varepsilon_{i},\beta\in\varepsilon_{j})$. This results in the synchronization of CTCs within the same ensemble as shown by Fig. 3(b,e,h). Surprisingly, considering all-to-all coupled CTCs ($\Gamma_{\alpha,\beta}/\kappa=\Gamma/\kappa=0.35~{}\forall~{}(\alpha,\beta)$) leads to the synchronization of the CTCs only in one ensemble (here $\varepsilon_{1}$), while the CTCs in the other ensemble (here $\varepsilon_{2}$) exhibit unsynchronized dynamics as shown in Fig. 3(c,f,i). Such surprising patterns of partial synchronization in systems of coupled identical nonlinear oscillators were first observed in Kuramoto and Battogtokh (2002) and were called chimera states Abrams and Strogatz (2004); Zakharova (2020). First strides towards chimera states in quantum systems involved exploring one-dimensional arrays of quantum oscillators Bastidas et al. (2015) and periodically driven networks of spins Sakurai et al. (2021). Our work demonstrates a chimera state in a system of coupled identical CTCs. Here, the chimera state results from including the asymmetric subspaces into the dynamics of identical CTCs, allowing us to choose an initial state with different frequencies in different subspaces Ujjwal et al. (2017b). Increasing the coupling strength results in the synchronization of all CTCs, see sup .
Another example of an exotic synchronization regime that can be realized in coupled identical CTCs is cluster synchronization, where the system separates into two differently synchronized clusters. Such a state is shown in Fig. 4(b), where the group $\varepsilon_{1}$ (black) and the group $\varepsilon_{2}$ (red) of CTCs get synchronized separately at different frequencies. This results from the choice of the initial state shown in Fig. 4(a), where the frequency distribution of the ensemble $\varepsilon_{1}$ is the same as in Fig. 3(a), while it ranges from 0.75 to 0.8 for $\varepsilon_{2}$.
Characterization of different phases.— We now extend our analysis to investigate the effect of choosing different initial states and coupling strengths $\Gamma$ on the synchronization properties of $n$ coupled CTCs. These CTCs are initialized in different $m$ subspaces such that the frequency of CTCs in each subgroup $\varepsilon_{1,2}$ are sampled from Gaussian distributions with distinct mean frequencies $\bar{\omega}_{1,2}$ and standard deviations $\sigma_{1,2}$. We investigate the synchronization and dynamical properties as a function of coupling strength $\Gamma$ and detuning $\Delta=\bar{\omega}_{2}-\bar{\omega}_{1}$ for fixed values of $\sigma_{1,2}$, where $\Delta$ and $\sigma_{1,2}$ depend on the choice of initial states. The mean value of Pearson’s correlation $\bar{\mathcal{C}}=\sum_{\alpha<\beta}\mathcal{C}_{\alpha\beta}/\mathcal{N}$ will be used to quantify the amount of synchronization, where $\mathcal{N}=n(n-1)/2$. This measure takes the value $\bar{\mathcal{C}}=0$ for no synchronization, $\bar{\mathcal{C}}=1$ for complete synchronization, and $10<\bar{\mathcal{C}}<$ 1 in the regime of partial synchronization, such as the chimera and cluster-synchronized states.
The synchronization measure alone is not sufficient to understand the complex Arnold-tongue structure observed in Fig. 4(e). Therefore, we also investigate the largest Lyapunov exponent $\lambda_{L}$ (Fig. 4(f)), which measures the sensitivity of the dynamics of a nonlinear system to its initial conditions Datseris and Parlitz (2022); Skokos et al. (2016). A positive value of $\lambda_{L}$ corresponds to chaotic dynamics where two nearby trajectories diverge exponentially over time, whereas the system settles down to a stable fixed point for $\lambda_{L}<0$. Finally, $\lambda_{L}\approx 0$ represents limit cycles and quasi-periodic dynamics.
In Figs. 4(e,f), we characterize different regimes making use of both the synchronization measure $\bar{\mathcal{C}}$ and the maximum Lyapunov exponent $\lambda_{L}$. For a finite detuning $\Delta$ and small coupling strength $\Gamma$, coupled CTCs exhibit unsynchronized dynamics indicated by $\bar{\mathcal{C}}\approx\lambda_{L}\approx 0$. Increasing $\Gamma$ leads to a regime of partial synchronization with $\bar{\mathcal{C}}>0$ and $\lambda_{L}\approx 0$, which includes chimera and cluster synchronization states. These two states can be distinguished by the correlation matrix $\mathcal{C}$ (see sup for more details). Further increasing $\Gamma$ leads to $\lambda_{L}<0$ and $\bar{\mathcal{C}}\approx 1$, causing the system to settle down into a highly-correlated fixed point. This sudden termination of oscillations is termed oscillation death. Higher values of $\Gamma$ lead to a less correlated chaotic mean-field dynamics, which is characterized by $\lambda_{L}>0$ and $\bar{\mathcal{C}}<1$. The reduced synchronization measure results from the chaotic nature of the dynamics which makes the system weakly correlated, see Fig. 4(e). Finally, even larger values of the coupling strength lead to complete synchronization such that $\bar{\mathcal{C}}=1$. Thus, depending on the parameters and initial state configuration, the system features various exotic synchronization regimes as shown in Fig. 4(e,f).
@@
Conclusions.— In this work, we focused on the fundamental question of whether continuous time crystals in spin models only exist if initialized in the symmetric subspace. Our first main result is that the inclusion of asymmetric subspaces leads to time-crystal behavior even if the ratio of drive to dissipation strength $\Omega/\kappa$ is less than one. A second result is that taking into account asymmetric subspaces leads to multistability: depending on the parameter values, the system can exhibit a time crystal or melted phase for different initial states. For an ensemble of coupled CTCs, this leads to our third result, the existence of chimera states and cluster synchronization, which is not possible for CTCs realized in the symmetric subspace. Finally, our system can also exhibit chaotic mean-field dynamics and oscillations death. Our results provide new insights into the dynamics of time-crystal phases and point out strategies in the design of quantum networks that may lead to the observation of exotic synchronization phenomena.
An interesting future direction is to study the effect of weak permutational symmetry-breaking interactions, which allow the system to move between various spin sectors. Such interactions are inevitable in various experimental settings and will offer further valuable insights into the fundamental characteristics of time crystals.
@@ Acknowledgements.
Acknowledgments.— P.S. and S.V. thank R. Fazio for discussions. Numerical calculations were performed using QuTiP Johansson et al. (2012, 2013) and PIQS Shammah et al. (2018). M.H. was supported by [MEXT Quantum Leap Flagship Program] Grants No. JPMXS0118067285 and No. JPMXS0120319794. P.S. and C.B. acknowledge financial support from the Swiss National Science Foundation individual grant (grant no. 200020 200481). S.V. acknowledges support from DST-QUEST grant number DST/ICPS/QuST/Theme-4/2019. While preparing the manuscript, we became aware of related work Iemini et al. (2024).
Appendix A Mean-field and fixed-point analysis of a single continuous time crystal
In this section, we extend the mean-field analysis of continuous time crystals (CTCs) outside the symmetric subspace discussed in the main text. The mean-field dynamics of the system described by Eq. (1) of the main text is governed by the following set of coupled differential equations
$\dot{m}_{x}$ $=\kappa m_{x}m_{z},$
$\dot{m}_{y}$ $=-\Omega m_{z}+\kappa m_{y}m_{z},$ (5)
$\dot{m}_{z}$ $=\Omega m_{y}-\kappa[m_{x}^{2}+m_{y}^{2}],$
where $m_{\alpha}=\langle S_{\alpha}\rangle/S$ are rescaled spin operators with $\alpha\in\{x,y,z\}$. For the symmetric Dicke subspace, $m$ is constrained to take the value $1$, which led to the breaking of time-translational symmetry at $\Omega/\kappa=1$ in earlier studies Iemini et al. (2018); Hajdušek et al. (2022). We will now consider the more general case $10\leq m<$ 1. The fixed points, $(m_{x}^{*},m_{y}^{*},m_{z}^{*})$, of the coupled differential Eqs. (B) for this case are given as $\mathcal{M}_{1}=(\pm m\sqrt{1-(m\kappa/\Omega)^{2}},m^{2}\kappa/\Omega,0)$ and $\mathcal{M}_{2}=(0,\Omega/\kappa,\pm\sqrt{m^{2}-(\Omega/\kappa)^{2}})$. The stability of these fixed points is given by the eigenvalues of the Jacobian $\mathcal{J}$, whose matrix elements are defined as $\mathcal{J}_{\alpha\beta}=d\dot{m}_{\alpha}/dm_{\beta}$. The eigenvalues of $\mathcal{J}$ corresponding to fixed points $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\{\Lambda_{1}\}=\{0,\pm\kappa\sqrt{m^{2}-(\Omega/\kappa)^{2}}\}$ and $\{\Lambda_{2}\}=\{0,\pm\kappa\sqrt{m^{2}-(\Omega/\kappa)^{2}},\pm\kappa\sqrt{m^{2}-(\Omega/\kappa)^{2}}\}$ respectively, where both the upper or both the lower signs have to be chosen together. Note that the dynamics of the system are now dependent on both the system parameters $\Omega/\kappa$ and rescaled total magnetization $m$ of the given spin-subspace. For $\Omega/\kappa<m$, $\mathcal{M}_{2}$ is the only physical state of the system since $\mathcal{M}_{1}$ has unphysical imaginary values for the expectation of spin components. The corresponding eigenvalues $\{\Lambda_{2}\}$ for the fixed point $\mathcal{M}_{2}$ are purely real, and hence it is a saddle fixed point of the system. Therefore, the system relaxes to a time-invariant state given by the fixed point $\mathcal{M}_{2}$ for initial states having $m>\Omega/\kappa$, and no time-crystal behavior is observed. For states with $m<\Omega/\kappa$, $\mathcal{M}_{1}$ is the only physical fixed point of the system. The eigenvalues $\{\Lambda_{1}\}$ for the fixed point $\mathcal{M}_{1}$ are purely imaginary, and hence, the corresponding fixed point is a center. Therefore, the system oscillates in time for initial states having $m<\Omega/\kappa$, and time-translational symmetry is broken even for $\Omega/\kappa<1$.
Using this fixed-point analysis, we can characterize the phase transition of the system as described by Fig. 1(a) of the main text. The system exhibits multistability for $\Omega/\kappa<1$. The two fixed points $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ have different basins of attraction: depending upon the initial state, the system can exhibit two different phases, namely, a melted and a time-crystal phase. For the Dicke subspace, this bistability does not exist, as depicted in Fig. 1(a) of the main text. The Dicke subspace corresponds to initial states having $m=1$, which shows that for $\Omega/\kappa<1$, the system always settles down to a saddle fixed point (dashed red line), and for $\Omega/\kappa>1$, it shows oscillatory behavior (solid red line). However, for $m<1$, both time-translation invariant and symmetry-broken phases coexist and are separated by the blue line corresponding to $m=\Omega/\kappa$. Initial states having $m>\Omega/\kappa$ on the left of $m=\Omega/\kappa$ (blue) line settle down to the saddle fixed point $\mathcal{M}_{2}$. The parameter region right to the $m=\Omega/\kappa$ (blue) line corresponds to $m<\Omega/\kappa$ and gives stable oscillations around the fixed point $\mathcal{M}_{1}$ leading to time-translational symmetry breaking.
To illustrate the phase-space portrait of the fixed points discussed above, we consider the following coordinate transformation $P=\tan^{-1}\left[m_{y}/m_{x}\right],~{}Q=m_{z}$ Iemini et al. (2018). In terms of the new co-ordinates $P$ and $Q$, the fixed points $\{P^{*},Q^{*}\}$ are given as $\mathcal{M}^{P,Q}_{1}=(\tan^{-1}[\pm 1/\sqrt{(\Omega/m\kappa)^{2}-1}],0)$ and $\mathcal{M}^{P,Q}_{2}=(\pm\pi/2,\pm\sqrt{m^{2}-(\Omega/\kappa)^{2}})$. The phase-space portrait of the CTC for different $m$ values is depicted in Fig. S1(a) for $\Omega/\kappa=0.75$. Initial states having $m<\Omega/\kappa$ keep oscillating in the vicinity of the fixed point $\mathcal{M}^{P,Q}_{1}$. As soon as we choose an initial state with $m>\Omega/\kappa$, the corresponding trajectory in Fig. S1(a) settles down to the saddle fixed point $\mathcal{M}^{P,Q}_{2}$. For $\Omega/\kappa>1$, there exists no physical state of the system that can have $m>\Omega/\kappa$ since $m\leq 1$ for the rescaled spin operators. The stability of the oscillations around the fixed point $\mathcal{M}_{1}^{P,Q}$ can be verified by observing Fig. S1(a), where each limit cycle trajectory consists of around 4000 cycles of oscillations. These trajectories are obtained by the time evolution of the exact mean-field equations (B) (including non-linear terms) and show no deviation even at such long times.
Figure S1(b) depicts the time evolution of the master equation (Eq. (1) of the main text) with the initial state belonging to the $m=0.1$ subspace, with $\{m_{x},m_{y},m_{z}\}=\{0,0,0.1\}$ for $\Omega/\kappa=0.9$. As the system size increases, the oscillation in $m_{z}$ becomes persistent and approaches the mean-field limit. Any initial state belonging to subspaces with $m>0.9$ will decay and approach a time-invariant steady state irrespective of system size. Thus, the exact treatment confirms the mean-field analysis for $\Omega/\kappa<1$.
@@ Appendix A Mean-field and fixed-point analysis of a single continuous time crystal
Appendix B Mean-field dynamics of coupled continuous time crystals
@@
Here we discuss the mean-field dynamics of the coherently coupled CTCs described by Eq. (3) of the main text. These equations are formulated in terms of the rescaled spin operators $m^{\alpha}_{\beta}=\langle\hat{S}^{\alpha}_{\beta}\rangle/S$ with $\alpha:\{x,y,z\}$ and $\beta:\{A,B\}$, and are defined as follows,
$\frac{dm^{x}_{\alpha}}{dt}$ $=\kappa m^{x}_{\alpha}m^{z}_{\alpha}+\frac{\Gamma}{n}m^{z}_{\alpha}\sum_{\beta\neq\alpha}m^{y}_{\beta},$
$\frac{dm^{y}_{\alpha}}{dt}$ $=-\Omega_{\alpha}m^{z}_{\alpha}+\kappa m^{y}_{\alpha}m^{z}_{\alpha}-\frac{\Gamma}{n}m^{z}_{\alpha}\sum_{\beta\neq\alpha}m^{x}_{\beta},$ (6)
@@
$\frac{dm^{z}_{\alpha}}{dt}$ $=\Omega_{\alpha}m^{y}_{\alpha}-\kappa\left[(m^{x}_{\alpha})^{2}+(m^{y}_{\alpha})^{2}\right]+\frac{\Gamma}{n}\left[m^{y}_{\alpha}\sum_{\beta\neq\alpha}m^{x}_{\beta}-m^{x}_{\alpha}\sum_{\beta\neq\alpha}m^{y}_{\beta}\right].$
@@
Appendix C Chimera state and synchronization via seeding
@@
The initial state leading to the chimera state shown in Fig. 3 of the main text is chosen such that the absolute values of the corresponding Jacobian eigenvalues $\{|\Lambda_{1}^{\alpha}|;\alpha=1,2,\ldots 20\}$ for the fixed point $\mathcal{M}_{1}$ are uniformly sampled in the window from $0.2$ to $0.25$ for CTCs in ensemble $\varepsilon_{1}$ (represented by black color) and from $0.75$ to $0.85$ for CTCs in ensemble $\varepsilon_{2}$ (represented by red color). For the given initial state, we investigate the effect of increasing the coupling strength between the all-to-all coupled CTCs on the synchronization properties. For coupling strength $\Gamma/\kappa=0.5$, all the CTCs in ensemble $\varepsilon_{1}$ are synchronized while CTCs in ensemble $\varepsilon_{2}$ only show partial synchronization, as shown in Figs. S2(a,d). Increasing $\Gamma$ we see more complex dynamics (not shown here), but the system continues to exhibit a chimera state. At $\Gamma/\kappa\approx 0.605$, a transition occurs: the strength of the coupling leads to the melting of the CTCs of ensemble $\varepsilon_{1}$ as depicted in Figs. S2(b,e). Further increasing the coupling strength leads to the synchronization of the CTCs in ensemble $\varepsilon_{2}$ which seeds oscillations to ensemble $\varepsilon_{1}$. As a result, at coupling strength $\Gamma/\kappa=1.2$, all the CTCs oscillate with a common frequency as shown in Figs. S2(c,f). Such a synchronization via seeding is consistent with what was observed in Hajdušek et al. (2022).
@@
Figure S2: Normalized Fourier transform $\mathcal{F}[m^{z}_{\alpha}]$(a-c) and time evolution of $m^{z}_{\alpha}$ (d-f) for coupling strength $\Gamma/\kappa=0.5$(a,d), $\Gamma/\kappa=0.605$(b,e) and $\Gamma/\kappa=1.2$(c,f). The frequencies of the uncoupled CTCs are uniformly sampled in the window from $0.2$ to $0.25$ in ensemble $\varepsilon_{1}$ and from $0.75$ to $0.85$ in ensemble $\varepsilon_{2}$.
Appendix D Cluster synchronization